OpenStudy (anonymous):

Limit of (x approaches infinity) of (sqrt (x^2 -3x+4) -x)

OpenStudy (anonymous):

Please can anyone help me out?

OpenStudy (anonymous):

\[\sqrt{(x^{2} -3x+4) -x}\] like this?

OpenStudy (anonymous):

No the -X is seperate. Limit(x>infinity) of (sqrt x^2 -3x+4) -x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt{x^{2}-3x+4}-x\] so question is, how big is square term compared to -x

OpenStudy (anonymous):

I dont know but the answer on the back of the book is -3/2

OpenStudy (anonymous):

do i have to divide everything by x^2?

OpenStudy (psymon):

Multiply top and bottom by the conjugate? \[\frac{ \sqrt{x^{2}-3x+4} - x }{ 1 }*\frac{ \sqrt{x^{2}-3x+4} + x }{ \sqrt{x^{2}-3x+4} + x }\] Then try and see what you can do.

OpenStudy (anonymous):

then i divide it by x which leaves me by -3/2. Thanks a lot!!!! :)

OpenStudy (psymon):

Yep yep :3